# Contracting with a Mechanism Designer

### v1

@@ Appendix C Missing Proofs
Now, we are ready to bound the upper bound of $\mathrm{PoDM}$. Define $\alpha^{*}=1-(\ln r)/r>0$. We know that $r(1-\alpha^{*})=\ln r$. Consequently,
$\mathrm{PoDM}(s,\ell)={}$ $\min_{\alpha\in[0,1]}\max_{n\in[1:\infty)}\frac{U_{\textsc{sb}}(n)}{U(\alpha,n)}$
[…]
@@
As a result, we finally derive that $\mathrm{PoA}(s,\ell)\geq\mathrm{PoDM}(s,\ell)\geq\frac{1}{\varepsilon}$. ∎
C.8 Proof of Lemma 7.4
See 7.4
Proof.
In this proof, we specially use $v_{\max}^{(n)}$ to denote the first-order statistic out of $n$ i.i.d variables drawing from distribution $F$. Let $\beta=\left\lceil\frac{\ell}{s}\right\rceil$. First, by definition, we have
$1-F^{\beta n}(z)\overset{(a)}{=}{}$ $\operatorname*{\mathrm{Pr}}\!\left[{v_{\max}^{(\beta n)}\geq z}\right]$
$\overset{(b)}{\leq}{}$ $\sum_{k=0}^{\beta-1}\operatorname*{\mathrm{Pr}}\!\left[{\max_{i\in[kn+1:(k+1)n]}\{v_{i}\}\geq z}\right]$
=$\overset{(c)}{=}{}$ $\sum_{k=0}^{\beta-1}\operatorname*{\mathrm{Pr}}\!\left[{v_{\max}^{(n)}\geq z}\right]$
=$\overset{(d)}{=}{}$ $\beta\cdot\left(1-F^{n}(z)\right),$
where equalities (a) and (d) hold since $F^{n}$ is the distribution of $v_{\max}^{(n)}$. Inequality (b) holds by partitioning the interval $[1:\beta n]$ into $\beta$ equal-sized subintervals, each of which has a size of $n$ and equality (c) holds since the variable $v_{i}$ is i.i.d. such that
$\operatorname*{\mathrm{Pr}}\!\left[{\max_{i\in[kn+1:(k+1)n]}v_{i}\geq z}\right]=\operatorname*{\mathrm{Pr}}\!\left[{v_{\max}^{(n)}\geq z}\right].$
Therefore, we finally obtain that
$\beta\cdot\left(1-F^{n}(z)\right)\geq 1-F^{\beta n}(z)\overset{(e)}{\geq}1-F^{\ell}(z),$
where inequality (e) holds since $\beta n\geq b$ and $1-F^{n}(z)$ increases with the increase of $n$. ∎
C.9 Proof of Lemma 7.5
See 7.5
The proof of Lemma 7.5 relies on the following technical lemmas.
Lemma C.3 (Babaioff et al., 2017, Lemma 5.3).
Let $F$ be a positive MHR distribution. Let $\mu_{n}$ be the except value of the first-order static out of $n$ i.i.d. variables drawing from distribution $F$. For all $s\leq\ell$, it holds
Lemma C.4.
Proof.
Recall that $\bar{\mu}_{n}$ is the excepted value of the first-order static out of $n$ i.i.d. variables drawing from distribution $\bar{F}$. We also let $\bar{\mu}_{0}=0$. Define $t=1-F(0)$ and $B(n,k)=\binom{n}{k}t^{k}(1-t)^{n-k}$. The first-best benchmark $U_{\textsc{fb}}(n)$ can be represented by $\bar{\mu}_{i}$ for all $i\in[n]$.
[…]
$U_{\textsc{fb}}(n)={}$ $\sum_{i=0}^{n}B(s,i)\mathcal{H}_{i}$
[…]
$=$ $\int_{0}^{1}\frac{(t-(1-t)z+1-t)^{n}-1}{z}\cdot{\mathrm{d}}z$
$=$ $\int_{0}^{1}\frac{(1-(1-t)z)^{n}-1}{z}\cdot{\mathrm{d}}z.$
$U_{\textsc{fb}}(n)=$ $\int_{0}^{1}\frac{(1-(1-t)z)^{n}-1}{z}\cdot{\mathrm{d}}z.$
which completes the proof. ∎
Lemma C.5.
Define ^$\hat{\mathcal{H}}_{0}=0$ and
^$\hat{\mathcal{H}}_{n}=\int_{0}^{1}\frac{1-(tz+1-t)^{n}}{1-z}{\mathrm{d}}z$
^$\frac{\hat{\mathcal{H}}_{s}}{\hat{\mathcal{H}}_{\ell}}\geq\frac{\ln(\frac{1}{2}ts+1)}{\ln(\frac{1}{2}t\ell+1)}.$
Proof.
By induction, it is sufficient to show that
^$\frac{\hat{\mathcal{H}}_{n+1}}{\hat{\mathcal{H}}_{n}}\leq\frac{\ln(\delta t(n+1)+1)}{\ln(\delta tn+1)},$
where $\delta=\frac{1}{2}$. Letting $\Delta_{n}=\hat{\mathcal{H}}_{n+1}-\hat{\mathcal{H}}_{n}$, it is sufficient to show that
which is equivalent to showing
We rewrite the above condition as
Therefore, by induction, it suffices to show
for any integer $n\in\mathbb{N}^{+}$.
Define real function $\xi\colon z\mapsto z\ln\left(\frac{1}{z}+1\right)$. We know $\xi(\cdot)$ is increasing since
which is equivalent to the fact that
Finally, according to Lemma C.6, we obtain that
which gives us the desired conclusion. ∎
Lemma C.6.
Define ^$\hat{\mathcal{H}}_{0}=0$ and
^$\hat{\mathcal{H}}_{n}=\int_{0}^{1}\frac{1-(tz+1-t)^{n}}{1-z}\cdot{\mathrm{d}}z,$
^$\frac{\hat{\mathcal{H}}_{n}-\hat{\mathcal{H}}_{n-1}}{\hat{\mathcal{H}}_{n+1}-\hat{\mathcal{H}}_{n}}\geq\frac{\frac{1}{2}tn+1}{\frac{1}{2}t(n-1)+1}.$
Proof.
$\frac{\Delta_{n-1}}{\Delta_{n}}\geq\frac{\delta tn+1}{\delta t(n-1)+1}.$
First, by the definition of ^$\hat{\mathcal{H}}_{n}$, we have
$\Delta_{n}={}$$\hat{\mathcal{H}}_{n+1}-\hat{\mathcal{H}}_{n}$
[…]
Observe that
$\Delta_{n-1}\cdot(\delta t(n-1)+1)-\Delta_{n}\cdot(\delta tn+1)$
[…]
This indicates that we only need to show the function
is non-negative when $z$ is a positive integer. Since $g_{t}(0)=0$, it is sufficient to show $g_{t}(\cdot)$ is weakly increasing for any fixed $t\in[0,1]$.
Next, notice that the first derivation of the function $g_{t}(\cdot)$ is
Clearly, we have $g^{\prime}_{1}(z)=2\delta z+1\geq 0$. Now, we assume that $t\in[0,1)$.
On the one hand, for any $t\in[0,1)$, we have
On the other hand, since $\ln(1-t)<0$ for $t\in[0,1)$, we have
which implies that the function $-g^{\prime}_{t}(z)/(1-t)^{z}$ is weakly decreasing. Combining the above two facts, we derive that for any $z\geq 0$, it holds $-g^{\prime}_{t}(z)/(1-t)^{z}\leq 0$, leading that $g^{\prime}_{t}(z)\geq 0$.
Now, we are ready to prove Lemma 7.5.
Proof of Lemma 7.5.
Let $\bar{F}$ be the conditional distribution conditioning on the event that the random value following $F$ is non-negative. Denote by $\bar{\mu}_{n}$ the expected value of the first-order statistic out of $n$ i.i.d. variables drawing from distribution $\bar{F}$. Define $t=1-F(0)$ and $B(n,k)=\binom{n}{k}t^{k}(1-t)^{n-k}$. According to Lemma C.4, we know
In addition, by using the Chernoff inequalities and Lemma C.3, we further obtain that
[…]
Recall that ^$\hat{\mathcal{H}}_{0}=0$ and
^$\hat{\mathcal{H}}_{n}=\int_{0}^{1}\frac{1-(tz+1-t)^{n}}{1-z}\cdot{\mathrm{d}}z,$
for any positive number $n\in\mathbb{N}^{+}$ and real $t\in[0,1]$. Consequentially, according to lemmas C.4 and C.5, it follows that
which completes the proof. ∎

### v2

@@ Appendix C Missing Proofs
Now, we are ready to bound the upper bound of $\mathrm{PoDM}$. Define $\alpha^{*}=1-(\ln r)/r>0$. Then, we know that $r(1-\alpha^{*})=\ln r$. By dividing the $[0,1]$ into two subintervals, we have
$\mathrm{PoDM}(1,\ell)={}$ $\min_{\alpha\in[0,1]}\max_{n\in[\ell]}\frac{U_{\textsc{sb}}(n)}{U(\alpha,n)}$
$\geq{}$ $\min_{\alpha\in[0,1]}\max\left\{\frac{U_{\textsc{sb}}(1)}{U(\alpha,1)},\frac{U_{\textsc{sb}}(\ell)}{U(\alpha,\ell)}\right\}$
$={}$ $\min\left\{\min_{\alpha\in[0,\alpha^{*}]}\max\left\{\frac{U_{\textsc{sb}}(1)}{U(\alpha,1)},\frac{U_{\textsc{sb}}(\ell)}{U(\alpha,\ell)}\right\},\min_{\alpha\in[\alpha^{*},1]}\max\left\{\frac{U_{\textsc{sb}}(1)}{U(\alpha,1)},\frac{U_{\textsc{sb}}(\ell)}{U(\alpha,\ell)}\right\}\right\}$
$\geq{}$ $\min\left\{\min_{\alpha\in[0,\alpha^{*}]}\frac{U_{\textsc{sb}}(1)}{U(\alpha,1)},\min_{\alpha\in[\alpha^{*},1]}\frac{U_{\textsc{sb}}(\ell)}{U(\alpha,\ell)}\right\}.$
@@
As a result, we finally derive that $\mathrm{PoA}(1,\ell)\geq\mathrm{PoDM}(1,\ell)\geq\frac{1}{\varepsilon}$. ∎